13

Given the function f (x) = x4 − 2x2 + 2,


the second derivative of f is

[a.      −10

[b.   10

[c.   12x2 − 4

[d.   2x(x − 5)


Question 14


Customers of a hardware store are willing to buy

Q = 80 − P 2

boxes of nails at P rand per box. Find the marginal revenue if the price per box is R3.

a.       R−53.00

b.       R71.00

c.       R53.00

d.       R−71.00


1
Expert's answer
2021-06-14T16:02:46-0400

13.

"f'(x)=4x^3-4x"

"f''(x)=12x^2-4"


Answer: [c].   12x2 − 4


14.

Total Revenue:

"TR=QP=80P-P^3=80\sqrt{80-Q}-(80-Q)^{3/2}"


 Marginal revenue:

"MR=\frac{dTR}{dQ}=-40/\sqrt{80-Q}+3\sqrt{80-Q}/2"


For the price R3:

"Q=80-9=71"

"MR=-40\/3+9\/2=-53\/6"